how can i get the % of numbers such as 1,000



-Thanks
Geo Gates

Answer by tutorcecilia(2144)

Use the formual for percents:
Part/Whole=Percent/100
.
Example: Find 12% of 1000
x/1000=12%/100
(100)(x)=(1000)(12) [Cross-multiply]
100x=12000 [Solve for x]
x=120
So, 12% of 1000 is 120
.
Example: What percent of 1000 is 350?
Part/Whole=Percent/100
350/1000=x%/100 [Plug-in the values]
(350)(100)=(1000)(x) [Cross-multiply]
35000=1000x
35%=x